OpenStudy (anonymous):

is this sum in simplest form? 3a/a+b + 3b/a+b = 3a+3b/a+b

OpenStudy (anonymous):

I assume you means 3a (a+b) + 3b/(a+b)

OpenStudy (anonymous):

*3a/(a+b)

OpenStudy (anonymous):

\[\frac{3a}{a+b}+\frac{3b}{a+b}=\frac{3a+3b}{a+b}\] is a start

OpenStudy (anonymous):

if so then nope because you can factor out a 3

OpenStudy (anonymous):

See my friend Satellite's equation above

OpenStudy (anonymous):

its true or false question

OpenStudy (anonymous):

Well can you factor a 3 out of the numerator?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

What is left after factoring out the 3?

OpenStudy (anonymous):

hey man is it in simplest form or not?

OpenStudy (anonymous):

what is the simplest form of \[\frac{3\times 5+3\times 7}{5+7}\]?

OpenStudy (anonymous):

JESUS CHRIST THE LORD

OpenStudy (anonymous):

Hey man ...sounds like you should get someone else to help...I am not into the spoon feeding game. Have fun.

OpenStudy (anonymous):

EricDEFINITELYNEEDSHELP